OpenStudy (anonymous):

How do i find all of the zeros for the equation: 3x^3-x^2-15x+5=0.

OpenStudy (anonymous):

You can try using the general formula for finding the roots of cubic functions, which is not trivial :( Instead, you can try to factor it :D

OpenStudy (anonymous):

so how would i factor a cubic number? i know how to factor a quadratic equation.

OpenStudy (anonymous):

i mean cubic equation.

OpenStudy (mathmate):

3x^3-x^2-15x+5 = 3x^3-x^2 -15x+5 =3x^3-x^2 -(15x-5) =x^2(3x-1) - 5(3x-1) =(x^2-5)(3x-1)

OpenStudy (anonymous):

Ok THANK YOU SOO MUCH!!!

OpenStudy (mathmate):

You're welcome! :)